OpenStudy (minecraftqueen):

-_- IF U CAN HELP ME ILL MEDAL If U = 6 feet, V = 8 feet, W = 3 feet, X = 7 feet, Y = 13 feet, and Z = 3 feet, what is the area of the object? 67 square feet 144 square feet 91 square feet 83 square feet

OpenStudy (minecraftqueen):

@Hero

OpenStudy (loser66):

pic?

OpenStudy (minecraftqueen):

@dan815

OpenStudy (minecraftqueen):

@dan815 https://app30.studyisland.com/userfiles/19373q5.gif

OpenStudy (minecraftqueen):

@dan815 https://app30.studyisland.com/userfiles/19373q5.gif

OpenStudy (loser66):

let me help @dan815

OpenStudy (minecraftqueen):

@dan815 https://app30.studyisland.com/userfiles/19373q5.gif

OpenStudy (loser66):

@dan815

OpenStudy (minecraftqueen):

@dan815 https://app30.studyisland.com/userfiles/19373q5.gif

OpenStudy (loser66):

@dan815

OpenStudy (minecraftqueen):

@loser66 sorry laptop lagg

OpenStudy (minecraftqueen):

@loser66 sorry laptop lagg

OpenStudy (minecraftqueen):

@loser66 sorry laptop lagg

OpenStudy (minecraftqueen):

@loser66 sorry laptop lagg

OpenStudy (loser66):



OpenStudy (loser66):



OpenStudy (loser66):

The area of the object = the area of the whole rectangle - area I - area II. And you know that the area of the whole rectangle = 13 * 11 the area of I = 3*4 the area of II = 8*6 You calculate your own problem.